OpenStudy (anonymous):

@Ankh A(n)=-2x2 A.-8;-64;-1,024 B. 1; -64;-16,384 C. -2;-16;-256 D. -4;-32;-512

OpenStudy (zzr0ck3r):

is that an x or a "times" what is the question?

OpenStudy (anonymous):

times and find the first, fourth, and eighth terms of the sequence

OpenStudy (anonymous):

@zzr0ck3r

OpenStudy (anonymous):

\[a_n=-2 \times 2\] We know that the equation for a geometric progression is: \[a_n = a_1 \times q^{n-1}\] So for -2x2 to occur n must be 1, since: \[x^1 = x\] So just from this we know that: \[a_1 = -2\] And: \[q = 2\]

OpenStudy (anonymous):

(correction) So for -2x2 to occur n must be 2

OpenStudy (anonymous):

We can plug in to test that the 4th and 8th terms of sequence apply and that the answer is indeed C. \[a_4 = -2 \times 2^{4-1} = -16\] \[a_8 = -2 \times 2^{8-1} = -256\]

OpenStudy (anonymous):

But be careful because: \[a \times b = b \times a\] So check if it isn't: \[a_n = 2^{n-1} \times -2\] That way a1 = 2, and we have no answer that starts with 2, so C can safely be chosen as the right one.